OpenStudy (anonymous):

hey plzz solve this for me m^2 +4 = 4m @mathmale

OpenStudy (xapproachesinfinity):

subtract 4m from both sides you will have quadratic equation you should know how to do it?

OpenStudy (anonymous):

cud u juts do and show me sir..i am confuseddd

OpenStudy (xapproachesinfinity):

\(\huge m^2-4m+4=0\)

OpenStudy (anonymous):

now???

OpenStudy (anonymous):

@xapproachesinfinity now what shud i do

OpenStudy (xapproachesinfinity):

what is that look like to you?

OpenStudy (xapproachesinfinity):

\(\large (a-b)^2=a^2-2ab+b^2\)

OpenStudy (anonymous):

but i dont know how to apply i trie dit still didnt get it sir

OpenStudy (xapproachesinfinity):

well a=m and b=2 can you see it now?

OpenStudy (xapproachesinfinity):

\(\huge m^2-4m+4=m^2-2\times2m+2^2\)

OpenStudy (anonymous):

sir just solvfe it plzz i forogt the method

OpenStudy (xapproachesinfinity):

well we cannot give answers! i can guide there! if you still stuck, this means you shouldn't take this class yet!

OpenStudy (anonymous):

hmmm :( just show me how to solve it plzzz...cuz i have a set of questions like this

OpenStudy (xapproachesinfinity):

\(\huge m^2-4m+4=(m-2)^2=0\) the means that \(\huge m-2=0~ \\\huge m=2\) that's you solution